OpenStudy (anonymous):

cos^2 ( x) if x = pi/4 am i suppose to do (cos pi/4)^2 or cos (pi/4)^2

OpenStudy (anonymous):

the first one...

OpenStudy (anonymous):

ok thanks for clearing the confusion!

OpenStudy (anonymous):

\[\huge cos^2\frac {\pi}{4}=(cos\frac {\pi}{4})^2 \]